OpenStudy (anonymous):

A current of 6A is drawn from a 120V line. What power is being developed?How much energy in joule and in kwt is expended if the current is drawn steadily for one week.

OpenStudy (ash2326):

Power= Current * Voltage Power=6 A* 120 V=720 watts If the current is drawn for one week Time T= 7*24*3600 seconds= 604800 seconds Energy= Power*Time= 720*604800 Joules or 435,456,000 Joules In kwt= 435,456,000/1000= 435456 kwt

OpenStudy (shayaan_mustafa):

use power formula, P=VI P=6*120 P=720W then, convert time in seconds from week. t=7*24*3600=604800s E=P*t E=720*604800 E=435456000W then, convert energy in KW (move point three decimal towards left) E=435456KW Regards, Read my Profile. :D